S_{1}-S_{0} VIBRONIC SPECTRA OF BENZENE CLUSTERS REVISITED: II. THE TRIMER

T. IIMORI, Y. AOKI, and \underline{Y. OHSHIMA}, Department of Chemistry, Graduate School of Science, Kyoto University, Kitashirakawa-Oiwakecho, Sakyo-ku, Kyoto 606-8502, Japan.

As the second part of our recent reinvestigation on the electronic spectra of benzene clusters,^{a} we present the trimer vibronic system studied by two-color (2C) resonance enhanced two-photon ionization (R2PI) and UV-UV holeburing experiments. This band system is observed only in the dimer-ion channel even with 2C-R2PI because of extensive fragmentation after photoionization, and thus it has been incorrectly assigned to an isomeric form of the dimer.^{b} Detailed examination with mixed samples of C_{6}H_{6} and C_{6}D_{6} has revealed that the parent neutral has a single isomeric form for each isotopomer, (C_{6}H_{6})_{m}(C_{6}D_{6})_{n} with m + n = 3. This observation confirms the equivalency in three benzene sites, which is consistent with the most stable cyclic form predicted by a NEMO calculation.^{c} One of the intermolecular modes shows prominent Franck-Condon activity, implying a substantial conformational change via photoexcitation. Other details observed in the vibrionc spectra, e.g., exitonic splitting, will also be discussed.

aT. Iimori and Y. Ohshima, J. Chem. Phys. 114, 2867 (2001).

bW. Scherzer, O. Kratzschmar, H. L. Selzle, and E. W. Schlag, Z. Naturforsch. 47a, 1248 (1992).

cO. Engkvist, P. Hobza, H. L. Selzle, and E. W. Schlag, J. Chem. Phys. 110, 5758 (1999).